If you roll two fair six-sided dice, what is the probability that the sum is 4 or higher?
Answer #1

   1   2   3   4   5   6
1   (1, 1)   (1, 2)   (1, 3)   (1, 4)   (1, 5)   (1, 6)
2   (2, 1)   (2, 2)   (2, 3)   (2, 4)   (2, 5)   (2, 6)
3   (3, 1)   (3, 2)   (3, 3)   (3, 4)   (3, 5)   (3, 6)
4   (4, 1)   (4, 2)   (4, 3)   (4, 4)   (4, 5)   (4, 6)
5   (5, 1)   (5, 2)   (5, 3)   (5, 4)   (5, 5)   (5, 6)
6   (6, 1)   (6, 2)   (6, 3)   (6, 4)   (6, 5)   (6, 6)

There are total 36 outcomes
There are 33 outcomes where sum is 4 or higher

Hence probability = 33/36 = 11/12